# The power-spectrum tensor in steady-state systems and its role in quantum friction

### v1

@@ II The power-spectrum tensor in steady-state systems
Therefore, we can conclude that $\underline{J}(\omega)$ fulfills the crossing relation $\underline{J}(\omega)=\underline{J}^{*}(-\omega)=\underline{J}^{\sf T}(-\omega)$, highlighting that $\underline{J}(\tau)$ is real. The vanishing of $\underline{J}(\omega)$ provides yet another version of the (equilibrium) FDT while $\underline{J}(\omega)\neq 0$ gives a measure of the violation of the detailed balance within NESS [70, 71, 72, 73, 26]. Specifically, for $T=0$ and $\omega>0$ we obtain
@@
$J_{ij}(\tau)=$ $\frac{1}{4\pi}\langle\{\hat{d}_{i}(\tau),\hat{d}_{j}(0)\}\rangle$
$-\frac{1}{(2\pi)^{2}}\mathcal{P}\int_{-\infty}^{\infty}\mathrm{d}\tau^{\prime}\;\frac{\langle\mathrm{i}[\hat{d}_{i}(\tau^{\prime}),\hat{d}_{j}(0)]\rangle}{\tau^{\prime}-\tau}~{},$ (16)
@@
which specifically demonstrates how the conditions for the validity of the FDT imply the vanishing of $\underline{J}$. Indeed, in equilibrium, the correlations $C_{ij}(\tau)=\langle\hat{d}_{i}(\tau)\hat{d}_{j}(0)\rangle=\langle\{\hat{d}_{i}(\tau),\hat{d}_{j}(0)\}\rangle/2-\mathrm{i}\langle\mathrm{i}[\hat{d}_{i}(\tau^{\prime}),\hat{d}_{j}(0)]\rangle/2$ become analytical so that their real and imaginary part must fulfill the Kramer-Kronig relations [74], for which the terms on the r.h.s. of Eq. (II) cancel to zero. More precisely, we would like to recall that the correlation tensor $\underline{C}(\tau)$ in the equilibrium state at a generic temperature $T$ satisfies the following two additional properties: Firstly, the elements of the matrix $\underline{C}^{*}(\tau)$ are analytical functions within the stripe $0<\tau_{I}<\hbar\beta$ in the upper half part of the complex $\tau$-plane [75]; Secondly, in an equilibrium state $\underline{C}(\tau)$ fulfills the so-called Kubo-Martin-Schwinger condition, i.e. $\underline{C}^{*}(\tau)=\underline{C}(\tau-\mathrm{i}\hbar\beta)$  [76, 3, 75, 74, 55].
@@
The previous considerations illustrate that, in addition to possibly having $\bm{\Omega}\neq 0$, the system’s quantum nonequilibrium dynamics is in general characterized by $\underline{J}\neq 0$. Although both quantities are connected to the breaking of the detailed balance occurring at equilibrium, the underlying concepts they are addressing are, in general, not identical. As we will see below, there are indeed situations for which $\bm{\Omega}=0$ and $\underline{J}\neq 0$.
III Viscous motion through the electromagnetic vacuum
@@
As soon as the system becomes more concrete, its specific properties and symmetries allow us to draw further conclusions on its stochastic dynamics. In this section we illustrate this aspect within the framework of quantum electrodynamics, where the quantum fluctuations pervading the system are those connected with the absorptive, dissipative and dispersive processes characterizing light-matter interaction. As a specific example, we consider a electrically neutral particle (e.g. an atom) which moves in vacuum with nonrelativistic velocity along a prescribed trajectory. We further assume that the trajectory of its center of mass is classical (usually a good approximation for particles with relatively large masses or velocities) and parallel to the translationally invariant direction of an arbitrary arrangement of electrically neutral material bodies, see Fig. 1. We denote the direction of motion with the unit vector $\mathbf{n}$ and we use the vector $\mathbf{R}_{a}$ to identify the particle’s lateral position. The material bodies are comprised of linear, reciprocal and passive materials which we further assume to be nonmagnetic, dissipative, possibly spatially dispersive and in thermal equilibrium with the surrounding electromagnetic field.
@@
Figure 1: Schematic representation of a particle moving in vacuum close to an arrangement of different, macroscopic objects. The particle moves on a straight line with direction $\mathbf{n}$, which coincides with the direction of translational invariance of the arrangement of macroscopic bodies. In this setup, the quantum fluctuations of the electromagnetic field behave as a viscous medium and the particle’s motion is affected by a corresponding drag force. An external force acting on the particle’s center of mass (not shown) keeps the motion at constant velocity $v$.
@@
Despite the system being electrically neutral, when the particle moves along its trajectory, it interacts with the quantum fluctuations of the electromagnetic field which, by the presence of the material bodies, are modified relative to vacuum. As a result of this interaction a frictional force acts on the particle [77, 18, 53] (see also below) that tends bring the particle to rest with respect to the privileged reference frame defined by the arrangement of the material bodies. Consequently, an external force acting on the particle’s center of mass is required to maintain the particle’s motion. When an external drive and the frictional force balance, the system reaches a nonequilibrium steady-state characterized by the particle moving at constant speed $v$.
@@
The electromagnetic viscosity experienced by the particle [54] is a manifestation of the fluctuating and dissipative nature of the system. Due to the existence of quantum fluctuations, it has the peculiarity to exist even at zero temperature, where it is usually referred to as quantum friction. Within the standard framework, the stationary fluctuating dynamics of the particle is usually described by its three-dimensional electric dipole (vector-)operator $\hat{\mathbf{d}}(t)$. In the above-discussed nonequilibrium steady-state, the corresponding correlation tensor $\underline{C}(\tau,v)$ and the power-spectrum tensor $\underline{S}(\omega,v)$ explicitly depend on the particle velocity $v$ which thus plays the role of a control variable of our nonequilibrium configuration [33]. Naturally, $v=0$ indicates the equilibrium state. In general, these tensors also depend on the position $\mathbf{R}_{a}$ but in order to avoid notational clutter, in the subsequent discussions, we suppress the dependence on $\mathbf{R}_{a}$ unless explicitly required. Clearly, for any value of $v$ and $\mathbf{R}_{a}$ the power-spectrum tensor and the related quantities must fulfill the relations discussed in Sec. II.
@@
In the NESS, the particle’s dipole does not only viscously move with velocity $v$ but, due to the interaction with its quantum electromagnetic environment, it can also gain and loose angular momentum [25, 26, 78]. However, owing to the system’s symmetry and its intrinsic dissipative behavior, in the stationary state the particle cannot have a net angular momentum with the direction parallel to $\mathbf{n}$. As a physical consequence, the dipole cannot feature a net averaged rotation around the direction of motion and Eq. (4) takes the form
$\underline{S}(\omega,v)=\underline{K}(\omega,v)+\bm{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}~{},$ (17)
@@
Equation (17) can be used as starting point for discussing some general aspects about the tensorial structure of $\underline{S}(\omega,v)$ and its functional dependence on $v$. Indeed, given that for $v=0$ the system is in equilibrium and, therefore, the power spectrum is real and symmetric, we can already conclude that $\bm{\Omega}_{\perp}(\omega,0)=0$. Since the consituent materials are reciprocal, we have
$\underline{S}(\omega,v)=\underline{M}_{n}\underline{S}(\omega,-v)\underline{M}^{-1}_{n}~{}.$ (18)
Here, we denote with $\underline{M}_{n}=\underline{\mathds{1}}-2\mathbf{n}\mathbf{n}^{\sf T}$ the transformation that describes the mirror symmetry with respect to the plane orthogonal to $\mathbf{n}$ ($\underline{\mathds{1}}$ is the identity matrix). The change in sign of the velocity results from applying $\underline{M}_{n}$ to the velocity vector $\mathbf{v}=v\mathbf{n}$, i.e. $\underline{M}_{n}\mathbf{v}=-v\mathbf{n}$. However, the same transformation leaves $\mathbf{R}_{a}$ and all (proper) vectors parallel to the plane of symmetry unchanged. Further, we find it convenient to decompose the real matrix $\underline{K}$ as
$\underline{K}(\omega,v)=\underline{\Delta}(\omega,v)+\bm{\chi}(\omega,v)\cdot\mathbf{\underline{\Lambda}}~{},$ (19)
where $\underline{\Delta}$ is diagonal and $\bm{\chi}$ is a three-dimensional real-valued vector. The matrices $[\underline{\Lambda}_{k}]_{ij}=|\epsilon_{kij}|$ are symmetric, traceless and trace-orthogonal to $\underline{L}_{i}$ and $\underline{\Delta}$. In particular, due to their connection with the Gell-Mann matrices [79], the decomposition over all the previous matrices is unique.
If we denote with $\underline{\Lambda}_{n}=\mathbf{n}\cdot\mathbf{\underline{\Lambda}}$ the matrix that corresponds to the direction $\mathbf{n}$, we obtain that $\underline{\Lambda}_{n}$ and $\underline{\Delta}$ are invariant with respect to $\underline{M}_{n}$. For the remaining matrices, we find the transformations $\underline{M}_{n}\bm{\underline{\Lambda}}_{\perp}\underline{M}^{-1}_{n}=-\bm{\underline{\Lambda}}_{\perp}$ and $\underline{M}_{n}\mathbf{\underline{L}}_{\perp}\underline{M}^{-1}_{n}=-\mathbf{\underline{L}}_{\perp}$. Due to the uniqueness of the decomposition, we can thus conclude from Eq. (18) that $\underline{\Delta}(\omega,v)$ and $\chi_{n}(\omega,v)$ are even in the velocity $v$, while $\bm{\chi}_{\perp}(\omega,v)$ and $\bm{\Omega}_{\perp}(\omega,v)$ are odd functions of $v$. Taking into account that $\underline{L}^{\sf T}_{i}=-\underline{L}_{i}$ we can further write (see App. B)
$\underline{S}(\omega,-v)=\underline{S}^{\sf T}(\omega,v)-2\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}~{},$ (20)
@@
$\underline{\alpha}_{\Im}(\omega,-v)=\underline{\alpha}_{\Im}^{\sf T}(\omega,v)-2\pi\frac{\bm{\chi}_{\perp}(\omega,v)-\bm{\chi}_{\perp}(-\omega,v)}{\hbar}\cdot\mathbf{\underline{\Lambda}}_{\perp}~{}.$ (21)
An interesting special case for this relation occurs if the expression for the susceptibility of the quantum dipole operator is identical to that of its classical counterpart or does not depend on temperature. Under these circumstances, the last term of Eq. (21) must always coincide with its classical limit. Since we already know that, classically, $\underline{S}^{*}(-\omega)=\underline{S}(\omega)$, we directly infer that $\bm{\chi}_{\perp}(-\omega,v)=\bm{\chi}_{\perp}(\omega,v)$. Therefore, the last term of Eq. (21) vanishes, leading to the identity $\underline{\alpha}_{\Im}(\omega,-v)=\underline{\alpha}_{\Im}^{\sf T}(\omega,v)$. Due to its analyticity and the Kramers-Kronig relations, this means that when the susceptibility has the same form within the quantum and classical description, it must fulfill $\underline{\alpha}(\omega,-v)=\underline{\alpha}^{\sf T}(\omega,v)$. This is tantamount to stating that, in this case, the system fulfills Onsager–Casimir-like relations where the atom’s speed $v$ takes the role of the external bias parameter [80].
If the system exhibits an additional mirror-symmetry $\underline{M}_{t}$, where $\mathbf{t}$ is a unit vector orthogonal to $\mathbf{n}$, then we can show along the lines discussed above that $\bm{\chi}=\chi_{t}\mathbf{t}$ and $\bm{\Omega}_{\perp}=\Omega_{t}\mathbf{t}$ (see App. B). However, while the symmetry described by $\underline{M}_{n}$ is independent of the transverse location $\mathbf{R}_{a}$, the symmetry associated with $\underline{M}_{t}$ can, in general, only exist when $\mathbf{R}_{a}$ is located within (or parallel to) the symmetry plane. In other words, conclusions about $\underline{M}_{t}$ do depend on the transverse position of the particle. Similarly, if an additional mirror-symmetry along the direction $\mathbf{n}\times\mathbf{t}$ exists, this necessarily implies that, when the particle moves along the associated symmetry axis (located at the intersection of the two corresponding symmetry planes), the power-spectral tensor is diagonal, suppressing any possible rotation of the dipole operator  [25, 78, 54]. As mentioned above, although in this case the particle is characterized by a specific position $\mathbf{R}_{a}$, this is an example of nonequilibrium dynamics for which $\bm{\Omega}=0$. In addition, we would like to point out that the above reasoning essentially identifies a specific physical reference frame with one of the axes coinciding with $\mathbf{n}$, within which the power-spectrum tensor has a diagonal form. Indeed, while a unitary transformation that diagonalizes the power-spectrum tensor always exists, this transformation does, in general, not correspond to a physical rotation.
@@
As a consequence, we find $\underline{C}(\tau,v)=\underline{C}^{\sf T}(\tau,v)$ and from the general properties of the correlation function we obtain $C_{\sigma}(\tau,v)=C^{*}_{\sigma}(-\tau,v)$. The corresponding power-spectrum tensor $\underline{S}(\omega,v)=\mathbf{d}\mathbf{d}^{\sf T}\Sigma(\omega,v)$ is real and symmetric, implying that within this model $\bm{\Omega}$ vanishes exactly, irrespective of the particle’s position. From the physical point of view, this is equivalent to the inability of the dipole vector to rotate around its initial value, which is consistent with having the real vector $\mathbf{d}$ being time-independent. This means that, within this model, the particle cannot absorb or emit angular momentum  [25]. Using this model, we obtain from symmetry considerations that $\Sigma(\omega,v)$ must be an even function of $v$. We can see this most easily by considering the case where the dipole operator is oriented orthogonal to the direction of motion, i.e. $\mathbf{d}\equiv\mathbf{d}_{\perp}$. In this case, Eq. (18) in conjunction with fact that the matrix $\mathbf{d}_{\perp}\mathbf{d}_{\perp}^{\sf T}$ is invariant with respect to $\underline{M}_{n}$, we find that $\Sigma(\omega,v)=\Sigma(\omega,-v)$.
III.1 The power spectrum in quantum friction

### v2

@@ 2 The power-spectrum tensor in steady-state systems
Therefore, we can conclude that $\underline{J}(\omega)$ fulfills the crossing relation $\underline{J}(\omega)=\underline{J}^{*}(-\omega)=\underline{J}^{\sf T}(-\omega)$, highlighting that $\underline{J}(\tau)$ is real. The vanishing of $\underline{J}(\omega)$ can yet provide another version of the (equilibrium) FDT while $\underline{J}(\omega)\neq 0$ gives a measure of the violation of the detailed balance within NESS [64, 65, 66, 67, 23]. Specifically, for $T=0$ and $\omega>0$ we obtain
@@
$J_{ij}(\tau)=$ $\frac{1}{4\pi}\langle\{\hat{d}_{i}(\tau),\hat{d}_{j}(0)\}\rangle-\frac{1}{(2\pi)^{2}}\mathcal{P}\int_{-\infty}^{\infty}\mathrm{d}\tau^{\prime}\;\frac{\langle\mathrm{i}[\hat{d}_{i}(\tau^{\prime}),\hat{d}_{j}(0)]\rangle}{\tau^{\prime}-\tau}~,$ (17)
@@
which specifically demonstrates how the conditions for the validity of the FDT imply the vanishing of $\underline{J}$. Indeed, in equilibrium, the correlations $C_{ij}(\tau)=\langle\hat{d}_{i}(\tau)\hat{d}_{j}(0)\rangle=\langle\{\hat{d}_{i}(\tau),\hat{d}_{j}(0)\}\rangle/2-\mathrm{i}\langle\mathrm{i}[\hat{d}_{i}(\tau^{\prime}),\hat{d}_{j}(0)]\rangle/2$ become analytical so that their real and imaginary part must fulfill the Kramer-Kronig relations [68], for which the terms on the r.h.s. of Eq. (17) cancel to zero  [^3] More precisely, we would like to recall that the correlation tensor $\underline{C}(\tau)$, in the equilibrium state at a generic temperature $T$, satisfies the following two additional properties: The elements of the matrix $\underline{C}^{*}(\tau)$ are analytical functions within the stripe $0<\tau_{I}<\hbar\beta$ in the upper half part of the complex $\tau$-plane [69]; In addition, in an equilibrium state $\underline{C}(\tau)$ fulfills the Kubo-Martin-Schwinger condition, i.e. $\underline{C}^{*}(\tau)=\underline{C}(\tau-\mathrm{i}\hbar\beta)$  [70, 3, 68, 69, 47]..
@@
The considerations above illustrate that, in addition to possibly having $\boldsymbol{\Omega}\neq 0$, the system’s quantum nonequilibrium steady state is in general characterized by $\underline{J}\neq 0$. Although both quantities are connected to the breaking of the detailed balance occurring at equilibrium, the underlying concepts they are addressing are, in general, not identical. As we will see below, there are indeed situations for which $\boldsymbol{\Omega}=0$ and $\underline{J}\neq 0$.
3 Viscous motion through the electromagnetic vacuum
@@
As soon as the system’s physical realization becomes more specific, its properties and symmetries allow us to draw further conclusions on its stochastic dynamics. In this section we illustrate this aspect within the framework of quantum electrodynamics, where the quantum fluctuations pervading the system are those connected with the absorptive, dissipative and dispersive processes characterizing light-matter interaction. As a concrete application, we consider an electrically neutral particle (e.g. an atom) which can move in vacuum with nonrelativistic velocity along a prescribed trajectory. We further assume that the trajectory of its center of mass is classical (usually a good approximation for particles with relatively large masses or velocities) and parallel to the translationally invariant direction of an arbitrary arrangement of electrically neutral material bodies, see Fig. 2. We denote the direction of motion with the unit vector $\mathbf{n}$ and we use the vector $\mathbf{R}_{a}$ to identify the particle’s lateral position. The material bodies are comprised of linear, reciprocal and passive materials which we further assume to be nonmagnetic, dissipative, possibly spatially dispersive and, in absence of the particle, in thermal equilibrium with the surrounding radiation.
@@
Figure 2: Schematic representation of a particle moving in vacuum close to an arrangement of different, macroscopic objects. The particle moves on a straight line with direction $\mathbf{n}$, which coincides with the direction of translational invariance of the arrangement of macroscopic bodies. In this setup, the quantum fluctuations of the electromagnetic field behave as a viscous medium and the particle’s motion is affected by a corresponding drag force. An external force acting on the particle’s center of mass (not shown) keeps the motion at constant velocity $v$. The power generated by the external drive is dissipated in the electromagnetic plus dissipative-matter environment [71, 23].
@@
Despite the system being electrically neutral, when the particle moves along its trajectory, it interacts with the fluctuations of the electromagnetic field which, by the presence of the material bodies, are modified relative to vacuum. As a result of this interaction, a frictional force acts on the particle [15, 43, 45] (see also below) that tends to bring the particle to rest with respect to the privileged reference frame defined by the arrangement of the material bodies. Consequently, an external force acting on center of mass of the subsystem particle is required to maintain its motion within the quantum environment resulting from the interaction of dissipative matter and the electromagnetic radiation. When an external drive and the frictional force balance, the system reaches a nonequilibrium steady-state characterized by the particle moving at constant speed $v$. In this state the input power, due to the drive acting on the atom, is exactly dissipated in the environment [71, 23].
@@
The electromagnetic viscosity experienced by the particle [72] is a manifestation of the fluctuating and dissipative nature of its quantum environment. Due to the existence of quantum fluctuations, the drag has the peculiarity to exist even at zero temperature, where it is usually referred to as quantum friction. Within the standard framework, the stationary fluctuating dynamics of the particle is usually described by its three-dimensional electric dipole (vector-)operator $\hat{\mathbf{d}}(t)$. In the above-discussed nonequilibrium steady-state, the corresponding correlation tensor, $\underline{C}(\tau,v)$, and the power-spectrum tensor, $\underline{S}(\omega,v)$, explicitly depend on the particle velocity $v$ which thus plays the role of a control variable of our nonequilibrium configuration (naturally, $v=0$ indicates in this case the equilibrium state) [37]. In general, these tensors also depend on the lateral position $\mathbf{R}_{a}$ but in order to avoid notational clutter, in the subsequent discussions, we suppress the dependence on $\mathbf{R}_{a}$ unless explicitly required. Clearly, for any value of $v$ and $\mathbf{R}_{a}$ the power-spectrum tensor and the related quantities must fulfill the relations discussed in Sec. 2.
@@
In the NESS, the particle’s dipole does not only viscously move with velocity $v$ but, due to the interaction with its quantum electromagnetic environment, it can also gain and loose angular momentum [22, 23, 73]. However, owing to the entire system’s symmetry and its intrinsic dissipative behavior, in the stationary state the particle cannot have a net angular momentum component with the direction parallel to $\mathbf{n}$. As a physical consequence, the dipole cannot feature a net averaged rotation around the direction of motion and Eq. (4) takes the form
$\underline{S}(\omega,v)=\underline{K}(\omega,v)+\boldsymbol{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}~,$ (18)
@@
Equation (18) can be used as starting point for discussing some general aspects about the tensorial structure of $\underline{S}(\omega,v)$ and its functional dependence on $v$. Indeed, given that for $v=0$ our system is in equilibrium, reciprocal and, therefore, the power spectrum is real and symmetric, we can already conclude that $\boldsymbol{\Omega}_{\perp}(\omega,0)=0$. Since the consituent materials are reciprocal, we have
$\underline{S}(\omega,v)=\underline{M}_{n}\underline{S}(\omega,-v)\underline{M}^{-1}_{n}~,$ (19)
which is a consequence of the system’s mirror-symmetry. Here, we denote with $\underline{M}_{n}=\underline{\mathds{1}}-2\mathbf{n}\mathbf{n}^{\sf T}$ the parity transformation that describes the reflection with respect to the plane orthogonal to $\mathbf{n}$ ($\underline{\mathds{1}}$ is the identity matrix). The change in sign of the velocity results from applying $\underline{M}_{n}$ to the velocity vector $\mathbf{v}=v\mathbf{n}$, i.e. $\underline{M}_{n}\mathbf{v}=-v\mathbf{n}$. However, the same transformation leaves $\mathbf{R}_{a}$ and all (proper) vectors parallel to the plane of symmetry unchanged. Further, we find it convenient to decompose the real matrix $\underline{K}$ as
$\underline{K}(\omega,v)=\underline{\Delta}(\omega,v)+\boldsymbol{\chi}(\omega,v)\cdot\mathbf{\underline{\Lambda}}~,$ (20)
where $\underline{\Delta}$ is diagonal and $\boldsymbol{\chi}$ is a three-dimensional real-valued vector. The matrices $\underline{\Lambda}_{k}]_{ij}=|\epsilon_{kij}|$ are symmetric, traceless and trace-orthogonal to $\underline{L}_{i}$ and $\underline{\Delta}$. In particular, due to their connection with the Gell-Mann matrices [74], the decomposition over all the previous matrices is unique.
If we denote with $\underline{\Lambda}_{n}=\mathbf{n}\cdot\mathbf{\underline{\Lambda}}$ the matrix that corresponds to the direction $\mathbf{n}$, we obtain that $\underline{\Lambda}_{n}$ and $\underline{\Delta}$ are invariant with respect to $\underline{M}_{n}$. For the remaining matrices, we find the transformations $\underline{M}_{n}\boldsymbol{\underline{\Lambda}}_{\perp}\underline{M}^{-1}_{n}=-\boldsymbol{\underline{\Lambda}}_{\perp}$ and $\underline{M}_{n}\mathbf{\underline{L}}_{\perp}\underline{M}^{-1}_{n}=-\mathbf{\underline{L}}_{\perp}$. Due to the uniqueness of the decomposition, we can thus conclude from Eq. (19) that $\underline{\Delta}(\omega,v)$ and $\chi_{n}(\omega,v)$ are even in the velocity $v$, while $\boldsymbol{\chi}_{\perp}(\omega,v)$ and $\boldsymbol{\Omega}_{\perp}(\omega,v)$ are odd functions of $v$. Taking into account that $\underline{L}^{\sf T}_{i}=-\underline{L}_{i}$ we can further write (see App. B)
$\underline{S}(\omega,-v)=\underline{S}^{\sf T}(\omega,v)-2\boldsymbol{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}~,$ (21)
@@
$\underline{\alpha}_{\Im}(\omega,-v)=\underline{\alpha}_{\Im}^{\sf T}(\omega,v)-2\pi\frac{\boldsymbol{\chi}_{\perp}(\omega,v)-\boldsymbol{\chi}_{\perp}(-\omega,v)}{\hbar}\cdot\mathbf{\underline{\Lambda}}_{\perp}~.$ (22)
An interesting special case for this relation occurs if the expression for the susceptibility of the quantum dipole operator is identical to that of its classical counterpart or does not depend on temperature. Under these circumstances, the last term on the r.h.s. of Eq. (22) must always coincide with its classical limit. Since we already know that, classically, $\underline{S}^{*}(-\omega)=\underline{S}(\omega)$, we directly infer that $\boldsymbol{\chi}_{\perp}(-\omega,v)=\boldsymbol{\chi}_{\perp}(\omega,v)$. Therefore, the last term of Eq. (22) vanishes, leading to the identity $\underline{\alpha}_{\Im}(\omega,-v)=\underline{\alpha}_{\Im}^{\sf T}(\omega,v)$. Due to its analyticity and the Kramers-Kronig relations, this means that when the susceptibility has the same form within the quantum and classical description, it must fulfill $\underline{\alpha}(\omega,-v)=\underline{\alpha}^{\sf T}(\omega,v)$. This is tantamount to stating that, in this case, the system fulfills Onsager–Casimir-like relations where the atom’s speed $v$ takes the role of the external bias parameter [75].
If the system exhibits an additional mirror-symmetry with respect to the reflection $\underline{M}_{t}$, where $\mathbf{t}$ is a unit vector orthogonal to $\mathbf{n}$, then we can show along the lines discussed above that $\boldsymbol{\chi}=\chi_{t}\mathbf{t}$ and $\boldsymbol{\Omega}_{\perp}=\Omega_{t}\mathbf{t}$ (see App. B). However, while the symmetry described by $\underline{M}_{n}$ is independent of the transverse location $\mathbf{R}_{a}$, the symmetry associated with $\underline{M}_{t}$ can, in general, only exist when $\mathbf{R}_{a}$ is located within (or parallel to) the symmetry plane. In other words, conclusions about $\underline{M}_{t}$ do depend on the transverse position of the particle. Similarly, if an additional mirror-symmetry along the direction $\mathbf{n}\times\mathbf{t}$ exists, this necessarily implies that, when the particle moves along the associated symmetry axis (located at the intersection of the two corresponding symmetry planes), the power-spectral tensor is diagonal, suppressing any possible rotation of the dipole operator  [22, 73, 72]. As mentioned above, although in this case the particle is characterized by a specific position $\mathbf{R}_{a}$, this is an example of nonequilibrium dynamics for which $\boldsymbol{\Omega}=0$. In addition, we would like to point out that the above reasoning essentially identifies a specific physical reference frame with one of the axes coinciding with $\mathbf{n}$, within which the power-spectrum tensor has a diagonal form. Indeed, while a unitary transformation that diagonalizes the power-spectrum tensor always exists, this transformation does, in general, not correspond to a physical rotation.
@@
As a consequence, we find $\underline{C}(\tau,v)=\underline{C}^{\sf T}(\tau,v)$ and from the general properties of the correlation function we obtain $C_{\sigma}(\tau,v)=C^{*}_{\sigma}(-\tau,v)$. The corresponding power-spectrum tensor $\underline{S}(\omega,v)=\mathbf{d}\mathbf{d}^{\sf T}\Sigma(\omega,v)$ is real and symmetric, implying that within this model $\boldsymbol{\Omega}$ vanishes exactly, irrespective of the particle’s position. From the physical point of view, this is equivalent to the inability of the dipole vector to rotate around its initial value, which is consistent with having the real vector $\mathbf{d}$ being time-independent. This means that, within this model, the particle cannot absorb or emit angular momentum  [22]. Using this model, we obtain from symmetry considerations that $\Sigma(\omega,v)$ must be an even function of $v$. We can see this most easily by considering the case where the dipole operator is oriented orthogonal to the direction of motion, i.e. $\mathbf{d}\equiv\mathbf{d}_{\perp}$. In this case, Eq. (19) in conjunction with fact that the matrix $\mathbf{d}_{\perp}\mathbf{d}_{\perp}^{\sf T}$ is invariant with respect to $\underline{M}_{n}$, we find that $\Sigma(\omega,v)=\Sigma(\omega,-v)$.
3.1 The power spectrum in quantum friction